OpenStudy (anonymous):

integrate; 1/(x+1)x^2

OpenStudy (amistre64):

x^-2 -x^-1 + 1 .... --------------- x^2+x^3 ) 1 (1+x) ------ -x (-x-x^2) --------- x^2 \[\int x^{-2} -x^{-1} + 1 -x+x^2-x^3....\ dx\]

OpenStudy (anonymous):

i did it by partial fraction decomposition

OpenStudy (amistre64):

i think the ideal thing to do tho is to decompose it into separate fraction

OpenStudy (amistre64):

1 a b c ------- = ---- + ---- + ------- x^2(x+1) x x^2 x+1 1 = ax(x+1) + b(x+1) + cx^2 1 = ax(x+1) + b(x+1) + cx^2 1 1 1 = ax(x+1) + (x+1) + x^2 1 = a2(3) + (3) + 9 1 = a6 + 12 a = -11/6 if I did that right

OpenStudy (amistre64):

missed it somples; but a=-1 according to the wolf

OpenStudy (anonymous):

help

OpenStudy (anonymous):

yeaah i got -1 for a and 1 for both B and C